OpenStudy (anonymous):

(sinx)/(1-cosx) + (sinx)/(1+cosx) = 2 csc x Need step by step! Trig homework help

OpenStudy (campbell_st):

well use the same method you would for adding 1/2 and 1/3 you need a common denominator... and then cross multiply... it starts as \[\frac{\sin(x)(1 + \cos(x)) + \sin(x)(1 - \cos(x))}{(1 - \cos(x))(1+\cos(x))}\] the denominator is the difference of 2 squares... and a variation on a key trig identity. distribute in the numerator and then collect like terms... then things may become clearer hope it helps